OpenStudy (anonymous):

A biologist is comparing the growth of a population of flies per week to the number of flies a bullfrog will consume per week. She has devised an equation to solve for which day (x) the bullfrog would be able to eat the entire population. The equation is 2^x = 3x - 1. However, she has observed that the bullfrog cannot eat more than seven flies in one week. Explain to the biologist how she can solve this on a graph using a system of equations. Identify any possible constraints to the situation.

OpenStudy (misssunshinexxoxo):

Previous OS question http://openstudy.com/study#/updates/5372fefbe4b0c1175e755dc4

OpenStudy (anonymous):

I looked at that, and I still don't understand how they went from "2^x = 3x -1" to "y = _____" . I don't understand where the "y" came from.

OpenStudy (misssunshinexxoxo):

Y is the dependent variable

OpenStudy (anonymous):

I understand that, but how do they get "y" if it wasn't even in the equation in the first place?

OpenStudy (anonymous):

Could you help me solve it step by step please?

OpenStudy (misssunshinexxoxo):

y = 3x -1

OpenStudy (anonymous):

2^x = 3x - 1 so then 2^x = y but y also equals "3x - 1"?

OpenStudy (misssunshinexxoxo):

Exactly

OpenStudy (anonymous):

y = 3x -1 y = 2^x 2^x = 3x-1 y = y????

OpenStudy (anonymous):

@misssunshinexxoxo I still don't understand

OpenStudy (anonymous):

Never mind, I figured it out.